OpenStudy (anonymous):

Let f(x) = 3x, x<2 x+10, 27 Explain the process you would go through to find f(8)

OpenStudy (hoblos):

x=8>7 => f(x)= 2-x = 2-8 = -6

OpenStudy (anonymous):

Can you please explain why its is X>7 ? I am having trouble figuring out what to do with x+10, 2\[2lex \le7\]

OpenStudy (anonymous):

\[2 \le x \le 7\]

OpenStudy (hoblos):

f(x)= 3x for x<2 f(x) = x+10 for 2<x<7 f(x)= 2-x for x>7 here you need to find f(8) then you must replace x=8 and since 8>7 then you use the third equation f(x)=2-x

OpenStudy (anonymous):

Thank you so much for clearing that up!

OpenStudy (hoblos):

my pleasure XD